A data set about speed dating includes "like" ratings of male dates made by the female dates. The summary statistics are n=196, x̅=6.87, s=1.99. Use a 0.05 significance level to test the claim that the population mean of such ratings is less than 7.00. Assume that a simple random sample has been selected. Identify the null and alternative hypotheses, test statistic, P-value, and state the final conclusion that addresses the original claim.

What are the null and altemative hypotheses?


Homework: Chapter 8 HW3 Sa Score: 0.17 of 1 pt 6 of 12 (12 complete) Hw Score: 42.69%, 5.12 of 12 8.3.14-T Question Help A data set about speed dating includes Iike ratings of male dates made by the female dates. The summary statistics are n = 196, x 6.87, s-1·99. Use a 0.05 significance level to test the claim that the population mean of such ratings is less than 7.00. Assume that a simple random sample has been selected. Identify the null and alternative hypotheses, test statistic, P-value, and state the final conclusion that addresses the original claim. What are the null and alternative hypotheses? OA, Ho : μ < 7.00 ус Но: #7.00 OB. Ho: H- 7.00 H H>7.00 D. Но: #7.00 H1 : μ #7.00 μ > 7.00 H1:H7.00 Determine the test statistic. (Round to two decimal places as needed.) Enter your answer in the answer box and then click Check Answer Clear All Check Answer
Answer #1

A)

Hypothesis:

Ho : μ = 7.00

H_{1}:\mu <7.00

Test Statistics:

n=196

\bar{x} = 6.87

s=1.99

  z=\frac{\bar{x}-\mu }{\frac{\sigma }{\sqrt{n}}}

  =\frac{6.87-7}{0.1421}

  =-0.9148

P -value is 0.180

It is calculated using normal dist table.

Here 0.05 < 0.180.

We accept the null hypothesis.

The population mean of such ratings is equal to 7.00

B) Hypothesis:

H_{0}:\mu =0

H_{1}:\mu >0

Test Statistics:

n=36

\bar{x} = 0.1

\sigma =23.6

  z=\frac{\bar{x}-\mu }{\frac{\sigma }{\sqrt{n}}} = \frac{0.1-0}{3.933} = 0.025

P-value is 0.490

It is calculated using normal dist table.

Here 0.01 < 0.490

Hence we accept the null hypothesis.

This concludes that with garlic treatment the mean change in LDL cholestrol is equal to zero.